OpenStudy (anonymous):

(62/2-32/8)-9 and please show how you solved it

OpenStudy (anonymous):

(62/2-32/8)-9 =(31-4)-9 =27-9 =18

OpenStudy (anonymous):

and how would you write (9x5-24)/3

OpenStudy (anonymous):

Is that 9 times 5 ?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

(9*5-24)/3 =(45-24)/3 =21/3 =7

OpenStudy (anonymous):

ty